# Detecting and Quantifying Isolated Singularities over Discrete Valuation Rings

### v1

@@ 4 Isolated Singularities over a DVR
An alternative proof of Proposition 4.14, similar to the proof of Proposition 3.6, follows from Theorem 3.23 in [de2020zariski]. It tells us that given some prime ideal $p\in\mathfrak{q}\subset V[[\underline{x}]]$, then $\langle f\rangle+J_{\delta}\left(f\right)\subset\mathfrak{q}$ if and only if $f\in\mathfrak{q}^{2}$, and so we get that $\left(\frac{V[[\underline{x}]]}{\langle f\rangle}\right)_{\mathfrak{q}}$ is not regular.
@@
First, by Proposition 4.14 we have that $N=ord_{f}\left(p\right)$ is finite and well defined. Now, denoting $r=\operatorname{ord}\left(f\right)$, then from Lemma 2.9 we can conclude that $\langle\left(\partial_{1}\left(f\right)\right)^{p},\dots,\left(\partial_{n}\left(f\right)\right)^{p}\rangle\subset\mathfrak{m}_{V}^{p\left(r-1\right)}\subset\mathfrak{m}_{V}^{r-1}$. Yet, since we have that $p^{N}\in\langle f,\left(\partial_{1}f\right)^{p},\dots,\left(\partial_{n}f\right)^{p}\rangle$, then there exists some $a\in V[[\underline{x}]]$ such that $p^{N}+af\in\langle\left(\partial_{1}\left(f\right)\right)^{p},\dots,\left(\partial_{n}\left(f\right)\right)^{p}\rangle$. Therefore, as $f\in\mathfrak{m}_{V}^{r}\subset\mathfrak{m}_{V}^{r-1}$, we can conclude that $p^{N}\subset\mathfrak{m}_{V}^{r-1}+\langle f\rangle\subset\mathfrak{m}_{V}^{r-1}$, and the result follows since the order of $p\in\mathfrak{m}_{V}$ is $1$. ∎
@@
Proposition 4.18 tells us that both $\mu(\cdot,\delta)$ and $\tau(\cdot,\delta)$ are semi-continuous with respect to $\delta$. This can be viewed as an analogue of Theorem 2.6 in [greuel2007introduction], of Lemma A.13 in [greuel2016right], and of item 1 in Remark 2.32.
As mentioned in Remark 4.12, a problem with Proposition 4.14 is that the value of $\tau(f,\delta)$ depends on the choice of $\delta$. But, by Remark 4.3 in [hochster2021jacobian], we have that the value
@@
The following lemma is a general result about local rings, inspired by Theorem 1.1. in [liu2018milnor], that we would like to use in order to bound $\tau(f,\delta)$.
@@
Let $(R,\mathfrak{m})$ be a local ring, let $I$ be an $\mathfrak{m}$-primary ideal, and assume $f\in R$ satisfies $f^{N}\in I$. Then we have that
$\operatorname{length}_{R}\left(\frac{R}{I}\right)\leq N\cdot\operatorname{length}_{R}\left(\frac{R}{I+\langle f\rangle}\right).$
@@
and by applying length over $R$ to this sequence we get $\textup{length}_{R}\left(\ker\left(\varphi\right)\right)=\textup{length}_{R}\left(\frac{R}{I+\langle f\rangle}\right)$. Since $f^{N}\in I$ we can conclude that $\langle f^{N-1}\rangle\cdot M\subset\ker\left(\varphi\right)$, where $M=\frac{R}{I}$. Therefore, we have a sequence of inclusions
@@
as desired. ∎
Remark 4.22.
As in Theorem 1.1. of [liu2018milnor], we have an equality in Lemma 4.21 if and only if $\ker\left(\varphi\right)=\langle f^{N-1}\rangle$. This is true since we have an equality if and only if for every $i$ we have that $\ker\left(\varphi\right)\cap\langle f^{i}\rangle=\ker\left(\varphi\right)$, but $\langle f^{N-1}\rangle\subset\ker\left(\varphi\right)$.
@@
and we can conclude that $\tau^{\Delta}\left(x^{2}+p^{2}\right)=\frac{2}{p}$. Therefore, by setting $\delta_{N}\left(x\right)=\left(p^{N-1}+1-p^{2p-2}-px^{2}g\right)^{0.5}$ for any $N<\operatorname{ord}_{f}\left(p\right)$ (this square root exists since $p^{N-1}+1-p^{2p-2}-px^{2}g$ is a unit), we have that $J_{\delta_{N}}\left(x^{2}+p^{2}\right)=\langle x^{2}+p^{2},x^{p},p^{N}\rangle$, giving different values for $\tau(f,\delta)$. In addition, for every $\delta\left(x\right)\in\mathfrak{m}_{V}$ we have that $\mu(f,\delta)=1$ (which is clearly finite). But for $\delta\left(x\right)=\left(-1+p^{2p-2}-px^{2}g+x\right)^{0.5}$ we have that $J_{\delta}\left(f\right)=\langle x^{p},px\rangle$, and so $\frac{V[[x]]}{J_{\delta}\left(f\right)}$ is one dimensional and therefore must have infinite length.
We end this section with an analogue of Proposition 4.14 over a ramified DVR, i.e., a complete DVR $(V,\langle\pi\rangle,\kappa)$ of mixed characteristic $(0,p)$ such that $\kappa$ is algebraically closed and $p\in\langle\pi^{2}\rangle$, based upon the work of KC in [kc2024singular]. Unlike the unramified case, the ramified case is much easier and can be directly compared with $\tau_{V}$. In particular, we have a special derivation that allows us to "derive with respect to $\pi$" directly.
We recall the main results in [kc2024singular] that are relevent for our discussion (specifically, Corollary 2.4 and Corollary 2.13). Note that in their paper, all results are over the ring of polynomials, but the results are valid over the ring of power series, as it is the completion of the ring of power series. It is a ramified analogue of Theorem 4.8.
@@
Let $f\in\mathfrak{m}_{V}$ and let $\tilde{f}$ be as in Notation 3.1. Then we define $\partial_{\pi}(f)=\text{pr}(\partial_{y}(\tilde{f}))\in V[[\underline{x}]]$, where pr is as in Remark 3.2. More explicitly, if $f=\sum_{\alpha}u_{\alpha}\pi^{n^{\alpha}}\underline{x}^{\alpha}$ as in Notation 3.1 then $\partial_{\pi}(f)=\sum_{\alpha}u_{\alpha}n_{\alpha}\pi^{n_{\alpha}-1}\underline{x}^{\alpha}$. Since $\partial_{\pi}(f)\,\operatorname{mod}\,\pi=\frac{\partial}{\partial\pi}(f)$, we can conclude that $\tau^{\pi}(f)=\operatorname{length}_{V[[\underline{x}]]}\left(\frac{V[[\underline{x}]]}{J(f)+\langle\pi,f,\partial_{\pi}(f)\rangle}\right)$. Thus, the result follows from the third item of Proposition 3.3. ∎
@@
    The proof of Proposition 4.28 shows us that we can view $\tau^{\pi}$ in fact as a ramified analogue of $\tau^{\Delta}$, since they both can be thought of as the co-length of $\langle f\rangle+J(f)$ plus the uniformizer and the respective "derivative with respect to the uniformizer".
@@
    Note that Proposition 4.28 gives us that if $f$ defines an isolated singularity then $\tau_{V}(f)<\infty$. This can be though of a generalization of Proposition 3.6 in the ramified case.
@@
Let $f_{1}=x^{2}+\pi^{p}$ and $f_{2}=x^{p}+\pi^{2}$. Observe that $f_{1}\sim f_{2}$. Yet we have that $\overline{f_{1}}=x^{2}$ and $\overline{f_{2}}=x^{p}$ and so $\frac{\partial}{\partial\pi}(f_{1})=\frac{\partial}{\partial\pi}(f_{2})=0$. Therefore $\tau^{\pi}(f_{1})=\dim_{\kappa}\left(\frac{\kappa[[x]]}{\langle x\rangle}\right)=1$ and $\tau^{\pi}(f_{2})=\dim_{\kappa}\left(\frac{\kappa[[x]]}{\langle x^{p}\rangle}\right)=p$.
Department of Mathematics, University of Utah, UT 84112
E-mail address: svoray@math.utah.edu

### v2

@@ 4 Isolated Singularities over a DVR
An alternative proof of Proposition 4.14, similar to the proof of Proposition 3.6, follows from Theorem 3.23 in [DGJ20]. It tells us that given some prime ideal $p\in\mathfrak{q}\subset V[[\underline{x}]]$, then $\langle f\rangle+J_{\delta}\left(f\right)\subset\mathfrak{q}$ if and only if $f\in\mathfrak{q}^{2}$, and so we get that $\left(\frac{V[[\underline{x}]]}{\langle f\rangle}\right)_{\mathfrak{q}}$ is not regular.
@@
First, by Proposition 4.14 we have that $N=ord_{f}\left(p\right)$ is finite and well defined. Now, denoting $r=\operatorname{ord}\left(f\right)$, then from Lemma 2.9 we can conclude that $\langle\left(\partial_{1}\left(f\right)\right)^{p},\dots,\left(\partial_{n}\left(f\right)\right)^{p}\rangle\subset\mathfrak{m}_{V}^{p\left(r-1\right)}\subset\mathfrak{m}_{V}^{r-1}$. Yet, as $p^{N}\in\langle f,\left(\partial_{1}f\right)^{p},\dots,\left(\partial_{n}f\right)^{p}\rangle$, there exists some $a\in V[[\underline{x}]]$ such that $p^{N}+af\in\langle\left(\partial_{1}\left(f\right)\right)^{p},\dots,\left(\partial_{n}\left(f\right)\right)^{p}\rangle$. Therefore, as $f\in\mathfrak{m}_{V}^{r}\subset\mathfrak{m}_{V}^{r-1}$, we can conclude that $p^{N}\subset\mathfrak{m}_{V}^{r-1}+\langle f\rangle\subset\mathfrak{m}_{V}^{r-1}$, and the result follows since the order of $p\in\mathfrak{m}_{V}$ is $1$. ∎
@@
From Proposition 4.18 we can conclude that both $\mu(\cdot,\delta)$ and $\tau(\cdot,\delta)$ are semi-continuous with respect to $\delta$. This can be viewed as an analogue of Theorem 2.6 in [GLS07], of Lemma A.13 in [GN16], and of item 1 in Remark 2.37.
As mentioned in Remark 4.12, a problem with Proposition 4.14 is that the value of $\tau(f,\delta)$ depends on the choice of $\delta$. But, by Remark 4.3 in [HJ21], we have that the value
@@
The following lemma is a general result about local rings, inspired by Theorem 1.1. in [LIU18] (and by Proposition 1 in [ALM23]), which we use to bound $\tau(f,\delta)$.
@@
Let $(R,\mathfrak{m})$ be a local ring, let $I$ be an $\mathfrak{m}$-primary ideal, and assume that $f\in R$ satisfies $f^{N}\in I$. Then we have that
$\operatorname{length}_{R}\left(\frac{R}{I}\right)\leq N\cdot\operatorname{length}_{R}\left(\frac{R}{I+\langle f\rangle}\right),$
with equality if and only if $(I\colon\langle f\rangle)=\langle f^{N-1}\rangle+I$.
@@
and by applying length over $R$ to this short exact sequence sequence we can conclude that $\textup{length}_{R}\left(\ker\left(\varphi\right)\right)=\textup{length}_{R}\left(\frac{R}{I+\langle f\rangle}\right)$. Since $f^{N}\in I$ we can conclude that $\langle f^{N-1}\rangle\cdot M\subset\ker\left(\varphi\right)$, where $M=\frac{R}{I}$. Therefore, we have a sequence of inclusions
@@
Note that we have equality if and only if $\ker(\varphi)=\frac{\langle f^{N-1}\rangle+I}{I}$. This is true since we have an equality if and only if $\ker\left(\varphi\right)\cap\frac{\langle f^{i}\rangle+I}{I}=\ker\left(\varphi\right)$ for every $i$ (i.e. $\ker(\varphi)\subset\frac{\langle f^{i}\rangle+I}{I}$), but $\frac{\langle f^{N-1}\rangle+I}{I}\subset\ker\left(\varphi\right)$. Yet, we have that $\ker(\varphi)=\{g+I\colon fg\in I\}=\frac{(I\colon\langle f\rangle)}{I}$, and the result follows. ∎
@@
and we can conclude that $\tau^{\Delta}\left(x^{2}+p^{2}\right)=\frac{2}{p}$. Therefore, for any for any $N<\operatorname{ord}_{f}\left(p\right)$ we can define $\delta_{N}\left(x\right)=\left(p^{N-1}+1-p^{2p-2}-px^{2}g\right)^{0.5}$ (this square root exists since $p^{N-1}+1-p^{2p-2}-px^{2}g$ is a unit), and we can conclude that $J_{\delta_{N}}\left(x^{2}+p^{2}\right)=\langle x^{2}+p^{2},x^{p},p^{N}\rangle$, giving different values for $\tau(f,\delta)$. In addition, for every $\delta\left(x\right)\in\mathfrak{m}_{V}$ we have that $\mu(f,\delta)=1$ (which is clearly finite). But for $\delta\left(x\right)=\left(-1+p^{2p-2}-px^{2}g+x\right)^{0.5}$ we have that $J_{\delta}\left(f\right)=\langle x^{p},px\rangle$, and so $\frac{V[[x]]}{J_{\delta}\left(f\right)}$ is one-dimensional and therefore must have infinite length.
We end this section with an analogue of Proposition 4.14 over a ramified DVR, i.e., a complete DVR $(V,\langle\pi\rangle,\kappa)$ of mixed characteristic $(0,p)$ such that $\kappa$ is algebraically closed and $p\in\langle\pi^{2}\rangle$, based upon the work of KC in [KC24]. Unlike the unramified case, the ramified case is much easier and can be directly compared with $\tau_{V}$. In particular, we have a special derivation that allows us to "derive with respect to $\pi$" directly.
We recall the main results in [KC24] that are relevent for our discussion (specifically, Corollary 2.4 and Corollary 2.13). Note that in his paper, all results are over the ring of polynomials, but the results are valid over the ring of power series, as it is the completion of the ring of power series. It is a ramified analogue of Theorem 4.8.
@@
Let $f\in\mathfrak{m}_{V}$ and let $\tilde{f}$ be as in Notation 3.1. Then we define $\partial_{\pi}(f)=\text{pr}(\partial_{y}(\tilde{f}))\in V[[\underline{x}]]$, where pr is as in Remark 3.2. More explicitly, if $f=\sum_{\alpha}u_{\alpha}\pi^{n^{\alpha}}\underline{x}^{\alpha}$ as in Notation 3.1 then $\partial_{\pi}(f)=\sum_{\alpha}u_{\alpha}n_{\alpha}\pi^{n_{\alpha}-1}\underline{x}^{\alpha}$. Since we have $\partial_{\pi}(f)\,\operatorname{mod}\,\pi=\frac{\partial}{\partial\pi}(f)$, we get that $\tau^{\pi}(f)=\operatorname{length}_{V[[\underline{x}]]}\left(\frac{V[[\underline{x}]]}{J(f)+\langle\pi,f,\partial_{\pi}(f)\rangle}\right)$. Thus, the result follows from the third item of Proposition 3.3. ∎
@@
    The proof of Proposition 4.27 tells us that we can view $\tau^{\pi}$ as a ramified analogue of $\tau^{\Delta}$, since they both can be thought of as the co-length of $\langle f\rangle+J(f)$ plus the uniformizer and the respective "derivative with respect to the uniformizer".
@@
    Proposition 4.27 gives us that if $f$ defines an isolated singularity then $\tau_{V}(f)<\infty$. This can be though of a generalization of Proposition 3.6 in the ramified case.
@@
Let $f_{1}=x^{2}+\pi^{p}$ and $f_{2}=x^{p}+\pi^{2}$. Note that $f_{1}\sim f_{2}$. Yet we have that $\overline{f_{1}}=x^{2}$ and $\overline{f_{2}}=x^{p}$ and so $\frac{\partial}{\partial\pi}(f_{1})=\frac{\partial}{\partial\pi}(f_{2})=0$. Therefore $\tau^{\pi}(f_{1})=\dim_{\kappa}\left(\frac{\kappa[[x]]}{\langle x\rangle}\right)=1$ and $\tau^{\pi}(f_{2})=\dim_{\kappa}\left(\frac{\kappa[[x]]}{\langle x^{p}\rangle}\right)=p$.